briyamilet1:

Solve the equation the square root of the quantity x plus 3 plus 4 equals 5 for the variable. Show each step of your solution process.

briyamilet1:

anyone? o.O

justjm:

Is this your question? \(\sqrt{x+3}+4=5\)

briyamilet1:

yes

justjm:

\(\sqrt{x+3}+4=5\) By the subtractive property of equality \(\sqrt{x+3}=1\) By the converse of the square root property of equality \(\left(\sqrt{x+3}\right)^{2}=1^{2}\) \(x+3=1\) Now you can continue

briyamilet1:

oki ty